OpenStudy (anonymous):

Which rule will find the nth term of the arithmetic sequence -58, -65, -72, -86, …? A.) an = -65 − 7n B.) an = -65 + 7n C.) an = 51 − 7n D.) an = -51 + 7n E.) an = -51 − 7n

hartnn (hartnn):

the first term = a1 = -58 can you find the common difference ?

OpenStudy (anonymous):

I can't

hartnn (hartnn):

common difference d is the difference between the consecutive terms, so, d= -65-(-58) = ... ?

OpenStudy (anonymous):

-7?

hartnn (hartnn):

yup! d=-7 now just plug that in \(a_n=a_1+(n-1)d\) a1 = -58

OpenStudy (anonymous):

it is an A.P with c,d =-65-(-72)=72-65=7 find tn by the formula tn=a+(n-1)d ,where a is first term and d is c.d

OpenStudy (anonymous):

sorry c.d=-72-(-65)=-75+65=-7

OpenStudy (anonymous):

So what do I do now?

hartnn (hartnn):

plug in d=-7 , a1 = -58 in \(a_n=a_1+(n-1)d\)

OpenStudy (anonymous):

-66?

OpenStudy (anonymous):

tn=-58+(n-1)*-7 tn=-58-7n+7=-51-7n

hartnn (hartnn):

how ? whats -58 -(-7) =... ?

OpenStudy (anonymous):

-51. And sorry I typed he equation wrong slightly.

OpenStudy (anonymous):

So it's E for the answer?

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

welcome ^_^